drug concentration: 60mg/ml, 20 ml vial

dose : 300mcg/kg/min

patient weight: 60 kg

duration of infusion: 5 hours

calculate the volume of the drug needed for the infusion, the dosing volume may be rounded to the nearest whole number. How many vials should be used for the patient? Please show work.
Answer #1

1ml=60 mg

1 vial =20 ml = 20* 60 =1200 mg

dose = 300 mcg/kg/min = 300*60 *60=1080000 mcg/hour mcg to mg = divided by 1000

= 1080000/1000=1080 mg/hour

dose should be administered for 5 hours = 1080* 5=5400 mg

volume in ml=5400/ 60 =90 ml (5400 mg/ 60 mg)

number of vials used = 90 ml/ 20 ml= 4.5